OpenStudy (anonymous):

−4 √x-3 =12 solve for x and identify if it is an extraneous solution

OpenStudy (jiteshmeghwal9):

\[-4\sqrt{x-3}=12\]by dividing both sides by -4. we have\[\sqrt{x-3}=-3\]now square both sides\[x-3=9\]now it is a linear equation in one variable. i think u can solve it now :)